# Non-conformally Einstein instantons in Conformal Gravity with and without nonlinear matter fields

### v1

@@ V Obstructions to conformally (anti-)self-dual Einstein spaces
This expression is nonvanishing as long as $b\neq 0$, even in the limit $\ell\to\infty$. In fact, switching off the integration constant $b$ also forces $c=0$, according to Eq. (14), thereby recovering the Kerr-NUT-AdS solution of General Relativity. Additionally, when evaluated on the one-parameter extension of the latter, the tensor $\mathcal{T}_{\mu\nu}$ in Eq. (17b) can be expressed as
@@
is a monotonically decreasing function of the radial coordinate, whose falloff towards the asymptotic boundary is $H\sim\mathcal{O}(r^{-2})$. For arbitrary values of the integration constants, this scalar is nonvanishing, even in the limit $\ell\to\infty$ and $b\neq 0$. Therefore, we conclude that the Kerr-NUT-AdS instanton of Conformal Gravity is not conformally related to its Einsteinian counterpart, as long as $b\neq 0$. This result extends the non-conformally Einstein metrics of Conformal Gravity studied in Ref. Liu et al. (2013). Even more, as the one-parameter extension of the Taub-NUT-AdS solution is continuously connected to that studied here in the limit $a\to 0$, one can conclude from the Dunajski-Tod theorem Dunajski and Tod (2014) that the solutions studied in Ref. Corral et al. (2021) are non-conformally Einstein as well.
@@ VI Coupling to nonlinear conformal matter
VII Non-linearly charged Gravitational instantons
Besides constant scalar fields, we are interested in gravitational instantons with self-gravitating conformally coupled scalars and ModMax fields, which introduce nonlinear effects while preserving conformal symmetry. Here, we construct a novel gravitational Taub-NUT instanton with self-gravitating, non-linear, and conformal matter. The solution can be regarded as a nonlinearly charged gravitational dyon, which is a gravitational solution with electric and magnetic mass. It represents a two-parameter extension of the original Taub-NUT in general relativity with a negative cosmological constant. Additionally, we also obtain a two-parameter extension of the Eguchi-Hanson space with a negative cosmological constant by introducing non-linear matter in Conformal Gravity.
@@ VII.1 Taub-NUT instanton with weakened AdS boundary conditions
This metric is asymptotically locally AdS, as can be seen by analyzing the asymptotic behavior of the Riemann tensor towards the asymptotic boundary. Nevertheless, it has a weakened AdS asymptotics due to the presence of the $b$-mode.
@@
Notice that, in the absence of matter fields, this solution reduces to that studied in Ref. Corral et al. (2021). In such a case, it is worth mentioning that this self-dual instanton is not conformally Einstein, since the constraints in Eq. (17) reduce to
@@
while $\mathcal{P}=0$. The first condition is nonvanishing as long as $b\neq 0$, and it remains valid even in the limit $\ell\to\infty$. In the case $b=0$, the constraint (42) further implies $c=0$, and one recovers the standard self-dual Taub-NUT metric, which is an Einstein space. In other words, the presence of the parameters $b$ and $c$ prevents the solution from being conformally related to an Einstein space. This analysis was not presented in Ref. Corral et al. (2021) and it provides a formal demonstration that the solution cannot be obtained from a conformal transformation of the Taub-NUT-AdS solution of Einstein gravity.
@@
$I_{\rm E}$ $=\frac{\pi^{2}}{72\nu}\left[(576q^{2}-3456)\nu+45+4c(c+3)-\frac{24n^{2}}{\ell^{2}}(2c-1152\nu+15)-\frac{144n^{4}}{\ell^{4}}(384\nu-5)\right],$ (45)
@@
Thus, if the restriction on the integration constants given in Eq. (51) is assumed, the solution reduces to the non-linearly charged dyonic Riegert metric dressed by scalar fields, whose specific form is given by
@@
Therefore, due to its nonvanishing temperature in the absence of scalar and Maxwell fields, the solution was interpreted as a thermalized vacuum in Ref. Lu et al. (2012). [^2] A similar behavior was found in unconventional conformal supergravity Alvarez et al. (2023) Nevertheless, the presence of non-linear conformal matter breaks the degeneracy and the partition function is no longer zero, as they contribute in a nontrivial way to the Euclidean on-shell action, which is given by
$I_{\rm E}=\frac{\pi\beta}{36\nu r_{+}}\left[12+c(12+c+192c\nu)\right]-\frac{2\pi\beta e^{\gamma}(p^{2}+q^{2})}{r_{+}}\,.$ (55)
@@ VII.2 Eguchi-Hanson
In Ref. Eguchi and Hanson (1978), Eguchi and Hanson constructed an asymptotically flat self-dual gravitational instanton in General Relativity, following a similar strategy of Belavin-Polyakov-Schwartz-Tyupkin in Yang-Mills theory Belavin et al. (1975). Then, generalizations of their solution were obtained by including the cosmological constant, Maxwell fields Eguchi and Hanson (1979), and also acceleration Chng et al. (2006). In higher-curvature gravity, extensions of the Eguchi-Hanson metric have been studied in Refs. Wong and Mann (2012); Hendi et al. (2012); Corral et al. (2021, 2022, 2025); Fenwick and Ghezelbash (2025). Additionally, it has been used as a seed metric to construct higher-dimensional gravitational solitons by oxidizing the original four-dimensional Eguchi-Hanson solution Clarkson and Mann (2006a, b); Durgut et al. (2023). In the presence of conformally-coupled scalars, a backreacted Eguchi-Hanson-like solution was found in Ref. Barrientos et al. (2022b). Here, we generalize the vacuum Eguchi-Hanson solution in Conformal Gravity obtained in Ref. Corral et al. (2021) by coupling the theory to spin-0 and spin-1 non-linear conformal matter.
@@
This metric is anti-self-dual, and it saturates the BPS bound of Conformal Gravity. In fact, notice that this space is not conformally Einstein, as can be verified from the Dunajski-Tod theorem by computing the conditions in Eq. (17), which lead to $\mathcal{P}=0$ and
@@
There are special limits where this solution becomes (anti-)self-dual. To verify this, we compute the following invariants
@@
$=4\pi\beta\left(\frac{c^{2}(5-576\nu)}{96\nu r_{+}^{4}}+q^{2}r_{+}^{-4e^{\gamma}}\right)\,$, (65)
@@ VIII Discussion
In this work, we studied non-conformally Einstein gravitational instantons in Conformal Gravity, both in vacuum and in the presence of nonlinear conformal matter that also preserves conformal symmetry. To this end, we focused on conformally coupled scalar fields and ModMax electrodynamics. We first revisited the Kerr-NUT-AdS extension constructed in Ref. Liu and Lu (2013), which possesses a vanishing Bach tensor. Since Conformal Gravity is finite for asymptotically locally AdS spaces, we computed the conserved charges of this solution using the Noether-Wald prescription Lee and Wald (1990); Iyer and Wald (1994); Wald (1993). In particular, we obtained its mass and angular momentum, which reduce to the corresponding expressions in general relativity in the limit $b\to 0$ and $c\to 0$ of the metric parameters $b,c$. However, other interesting limits are possible. After performing the continuation to Euclidean signature, we explored the curve in parameter space along which the solution becomes globally (anti)-self-dual. Using the Dunajski-Tod theorem Dunajski and Tod (2014), we showed that the one-parameter extension of the Kerr metric is not conformally Einstein. Its Euclidean on-shell Euclidean action was obtained, which turns out to be proportional to the Chern-Pontryagin index, thereby saturating a gravitational BPS bound. We also demonstrated that this (anti)-self-dual instanton dominates the path integral over the complete projective space $\mathbb{CP}^{2}$, previously studied by Strominger, Horowitz, and Perry in Ref. Strominger et al. (1984).
In the second part of the paper, we found gravitational instantons in Conformal Gravity with conformally-coupled scalars and ModMax fields. In particular, we presented the generalization of two well-known gravitational instantons of general relativity: the Taub-NUT-AdS and Eguchi-Hanson metrics. For both cases, their temperature, the curve in parameter space where they become (anti)-self-dual, their partition function, and nonlinear Abelian charges were obtained. Similarly to the Kerr-NUT-AdS extension, we used the Dunajski-Tod theorem Dunajski and Tod (2014) to show that the vacuum Taub-NUT-AdS instanton in Conformal Gravity is not conformally Einstein. Its Euclidean on-shell action was computed following the prescription in Ref. Anastasiou et al. (2023), where suitable boundary terms are added without breaking the conformal invariance of the action. The electric and magnetic charges associated with the ModMax fields, defined through a Gauss law and through the integral of the first Chern class, respectively, were also computed. We showed that these charges are directly related to the two integration constants appearing in the ModMax field configuration. Furthermore, we analyzed the static limit of the NUT charge $n\to 0$, in which the solution reduces to the massless limit of the Riegert metric Riegert (1984) dressed with nonlinear conformal matter.
@@ Acknowledgements.
The authors thank Adolfo Cisterna for useful discussions. C.C. thanks the National Technical University of Athens for its hospitality during the initial stage of this work. This work is partially supported by Agencia Nacional de Investigación y Desarrollo, Chile, via Fondecyt Regular grants 1240043, 1240048, 1251523, 1252053, and 1230112.

### v2

@@ V Obstructions to conformally (anti)self-dual Ricci-flat spaces
This expression is nonvanishing as long as $b\neq 0$, even in the limit $\ell\to\infty$. Therefore, the nonvanishing of Eq. (17) implies that the solution is not conformal to a Ricci-flat space. In fact, switching off the integration constant $b$ also forces $c=0$, according to Eq. (14), thereby recovering the Kerr-NUT-AdS solution of General Relativity. Additionally, when evaluated on the one-parameter extension of the latter, the tensor $\mathcal{T}_{\mu\nu}$ in Eq. (17b) can be expressed as
@@
is a monotonically decreasing function of the radial coordinate, whose falloff toward the asymptotic boundary is $H\sim\mathcal{O}(r^{-2})$. For arbitrary values of the integration constants, this scalar is nonvanishing, even in the limit $\ell\to\infty$ and $b\neq 0$. Therefore, we conclude that the Kerr-NUT-AdS instanton of Conformal Gravity is not conformally related to its Ricci-flat counterpart, as long as $b\neq 0$. This result extends the nonconformally Einstein metrics of Conformal Gravity studied in Ref. Liu et al. (2013). Even more, as the one-parameter extension of the Taub-NUT-AdS solution is continuously connected to that studied here in the limit $a\to 0$, one can conclude from the Dunajski-Tod theorem Dunajski and Tod (2014) that the solutions studied in Ref. Corral et al. (2021) are nonconformally Ricci-flat as well.
@@ VI Coupling to nonlinear conformal matter
VII Nonlinearly charged Gravitational instantons
Besides constant scalar fields, we are interested in gravitational instantons with self-gravitating conformally coupled scalars and ModMax fields, which introduce nonlinear effects while preserving conformal symmetry. Recently, the extension of the Riegert black hole with meronic fields was found in Conformal Gravity minimally coupled to Yang-Mills theory Flores-Alfonso (2025). Here, we construct a novel gravitational Taub-NUT instanton with self-gravitating, nonlinear, and conformal matter. The solution can be regarded as a nonlinearly charged gravitational dyon, which is a gravitational solution with electric and magnetic mass. It represents a two-parameter extension of the original Taub-NUT in general relativity with a negative cosmological constant. Additionally, we also obtain a two-parameter extension of the Eguchi-Hanson space with a negative cosmological constant by introducing nonlinear matter in Conformal Gravity.
@@ VII.1 Taub-NUT instanton with weakened AdS boundary conditions
This metric is asymptotically locally AdS, as can be seen by analyzing the asymptotic behavior of the Riemann tensor toward the asymptotic boundary. Nevertheless, it has a weakened AdS asymptotics due to the presence of the $b$-mode.
@@
Notice that, in the absence of matter fields, this solution reduces to that studied in Ref. Corral et al. (2021). In such a case, it is worth mentioning that this self-dual instanton is not conformally Ricci-flat either, since the constraints in Eq. (17) reduce to
@@
while $\mathcal{P}=0$. The first condition is nonvanishing as long as $b\neq 0$, and it remains valid even in the limit $\ell\to\infty$. In the case $b=0$, the constraint (42) further implies $c=0$, and one recovers the standard self-dual Taub-NUT metric, which is an Einstein space. In other words, the presence of the parameters $b$ and $c$ prevents the solution from being conformally related to a Ricci-flat space. This analysis was not presented in Ref. Corral et al. (2021), and it provides a formal demonstration that the solution cannot be obtained from a conformal transformation of the Taub-NUT-AdS solution of Einstein gravity.
@@
$I_{\rm E}$ $=\frac{\pi^{2}}{72\nu}\left[(576q^{2}-3456\alpha)\nu+45+4c(c+3)-\frac{24n^{2}}{\ell^{2}}(2c-1152\nu\alpha+15)-\frac{144n^{4}}{\ell^{4}}(384\nu\alpha-5)\right],$ (45)
@@
Thus, if the restriction on the integration constants given in Eq. (51) is assumed, the solution reduces to the nonlinearly charged dyonic Riegert metric dressed by scalar fields, whose specific form is given by
@@
Therefore, due to its nonvanishing temperature in the absence of scalar and Maxwell fields, the solution was interpreted as a thermalized vacuum in Ref. Lu et al. (2012). [^3] A similar behavior was found in unconventional conformal supergravity Alvarez et al. (2023) Nevertheless, the presence of nonlinear conformal matter breaks the degeneracy and the partition function is no longer zero, as they contribute in a nontrivial way to the Euclidean on-shell action, which is given by
$I_{\rm E}=\frac{\pi\beta}{36\nu r_{+}}\left[12+c(12+c+192c\nu\alpha)\right]-\frac{2\pi\beta e^{\gamma}(p^{2}+q^{2})}{r_{+}}\,.$ (55)
@@ VII.2 Eguchi-Hanson
In Ref. Eguchi and Hanson (1978), Eguchi and Hanson constructed an asymptotically flat self-dual gravitational instanton in General Relativity, following a similar strategy of Belavin-Polyakov-Schwartz-Tyupkin in Yang-Mills theory Belavin et al. (1975). Then, generalizations of their solution were obtained by including the cosmological constant, Maxwell fields Eguchi and Hanson (1979), and also acceleration Chng et al. (2006). In higher-curvature gravity, extensions of the Eguchi-Hanson metric have been studied in Refs. Wong and Mann (2012); Hendi et al. (2012); Corral et al. (2021, 2022, 2025); Fenwick and Ghezelbash (2025). Additionally, it has been used as a seed metric to construct higher-dimensional gravitational solitons by oxidizing the original four-dimensional Eguchi-Hanson solution Clarkson and Mann (2006a, b); Durgut et al. (2023). In the presence of conformally-coupled scalars, a backreacted Eguchi-Hanson-like solution was found in Ref. Barrientos et al. (2022b). Here, we generalize the vacuum Eguchi-Hanson solution in Conformal Gravity obtained in Ref. Corral et al. (2021) by coupling the theory to spin-0 and spin-1 nonlinear conformal matter.
@@
This metric is anti-self-dual, and it saturates the BPS bound of Conformal Gravity. In fact, notice that this space is not conformally Ricci-flat, as can be verified from the Dunajski-Tod theorem by computing the conditions in Eq. (17), which lead to $\mathcal{P}=0$ and
@@
There are special limits where this solution becomes (anti)self-dual. To verify this, we compute the following invariants
@@
$=4\pi\beta\left(\frac{c^{2}(5-576\nu\alpha)}{96\nu r_{+}^{4}}+q^{2}r_{+}^{-4e^{\gamma}}\right)\,$, (65)
@@ VIII Discussion
In this work, we studied nonconformally Ricci-flat gravitational instantons in Conformal Gravity, both in vacuum and in the presence of nonlinear conformal matter that also preserves conformal symmetry. To this end, we focused on conformally coupled scalar fields and ModMax electrodynamics. We first revisited the Kerr-NUT-AdS extension constructed in Ref. Liu and Lu (2013), which possesses a vanishing Bach tensor. Since Conformal Gravity is finite for asymptotically locally AdS spaces, we computed the conserved charges of this solution using the Noether-Wald prescription Lee and Wald (1990); Iyer and Wald (1994); Wald (1993). In particular, we obtained its mass and angular momentum, which reduce to the corresponding expressions in general relativity in the limit $b\to 0$ and $c\to 0$ of the metric parameters $b,c$. However, other interesting limits are possible. After performing the continuation to Euclidean signature, we explored the curve in parameter space along which the solution becomes globally (anti)-self-dual. Using the Dunajski-Tod theorem Dunajski and Tod (2014), we showed that the one-parameter extension of the Kerr metric is not conformally Ricci-flat. Its Euclidean on-shell Euclidean action was obtained, which turns out to be proportional to the Chern-Pontryagin index, thereby saturating a gravitational BPS bound. We also demonstrated that this (anti)-self-dual instanton dominates the path integral over the complete projective space $\mathbb{CP}^{2}$, previously studied by Strominger, Horowitz, and Perry in Ref. Strominger et al. (1984).
In the second part of the paper, we found gravitational instantons in Conformal Gravity with conformally-coupled scalars and ModMax fields. In particular, we presented the generalization of two well-known gravitational instantons of general relativity: the Taub-NUT-AdS and Eguchi-Hanson metrics. For both cases, their temperature, the curve in parameter space where they become (anti)-self-dual, their partition function, and nonlinear Abelian charges were obtained. Similarly to the Kerr-NUT-AdS extension, we used the Dunajski-Tod theorem Dunajski and Tod (2014) to show that the vacuum Taub-NUT-AdS instanton in Conformal Gravity is not conformally Ricci-flat. Its Euclidean on-shell action was computed following the prescription in Ref. Anastasiou et al. (2023), where suitable boundary terms are added without breaking the conformal invariance of the action. The electric and magnetic charges associated with the ModMax fields, defined through a Gauss law and through the integral of the first Chern class, respectively, were also computed. We showed that these charges are directly related to the two integration constants appearing in the ModMax field configuration. Furthermore, we analyzed the static limit of the NUT charge $n\to 0$, in which the solution reduces to the massless limit of the Riegert metric Riegert (1984) dressed with nonlinear conformal matter.
@@ Acknowledgements.
The authors thank Eloy Ayón-Beato, Adolfo Cisterna, and Daniel Flores-Alfonso for useful discussions. C.C. thanks the National Technical University of Athens for its hospitality during the initial stage of this work. This work is partially supported by Agencia Nacional de Investigación y Desarrollo, Chile, via Fondecyt Regular grants 1240043, 1240048, 1251523, 1252053, and 1230112.